# The unreasonable likelihood of being Origin of life, terraforming, and AI

### v1

@@ Abstract
An alternative possibility is that life arose via a sudden explosive-like transition — by some yet-unknown self-organization principle of collectivessness triumpying over reductionism, that bridged the gap between random chemistry and selective evolution, potentially circumventing the information bottleneck of early self-replication. This idea is not new. It is most well developed in the study of collectively autocatalytic sets, first introduced by Kauffman (1986) [56]. In such networks, each molecule is catalyzed by another member of the set, enabling system-wide replication without requiring any single molecule to self-replicate or the presence of a membrane-enclosed protocell. This framework offers a resolution to Eigen’s paradox: long, information-rich molecules require error correction to persist, but error-correcting machinery itself demands such complexity [57]. Autocatalytic sets enable increasing functional complexity before templated replication emerges. Mathematical models show that as chemical diversity increases, the probability of forming an autocatalytic set undergoes a sharp phase transition—a sudden onset of self-sustaining autocatalysis [58, 59]. This is analogous to a percolation transition in random graphs.
@@
Such transitions may also explain later leaps in complexity, including the emergence of the brain. Put simply, a brain functions as an internal model of the external world, trained by experience and used to predict future outcomes [60] [^1] This predictive capacity has been formalized in the Bayesian brain hypothesis [61], where internal priors are continuously updated in light of new data, approximating posterior beliefs that guide perception and action.. For example, one can model parameter adaptation in a toy brain using the inverse temperature $\beta$ in an Ising-like system, rather than connectivity or size, as the critical control parameter. In such models, evolutionary fitness peaks near a critical value of $\beta$ when environments are dynamic, heterogeneous, and unpredictable [60]. Though phase transitions technically apply to infinite systems, the benefits of sharp transitions often hold in large finite systems as well [60, 62, 63].
@@
In the spirit of this article, we now return to AI. Suppose that chemical reaction networks in a cell can be viewed as a kind of neural network. Then we may ask: how much information must be encoded in such a network to achieve life-like complexity? Recent work on recurrent neural chemical reaction networks (RNCRNs) has shown that artificial recurrent neural networks can be mapped onto chemical reaction networks [64]. These networks are capable of approximating arbitrary dynamical systems governed by ordinary differential equations. Unlike feedforward neural networks, which approximate static input–output mappings, RNCRNs model time-dependent behaviors such as switching in multi-stable systems or oscillations frequently observed in biology.
@@
Currently, such RNCRNs remain small ($<10$ species or nodes) [64], and it remains unclear how large a network must be to compute the full repertoire of cellular behavior. Still, we can try and estimate the required network size. The universal approximation theorem states that for any continuous real-valued function and any target error $\varepsilon>0$, there exists a neural network that approximates $f$ within $\varepsilon$ on a compact domain. Recent results give an upper bound on the number of neurons $M$ needed to achieve this approximation:
@@
Let us now take a different route and investigate how large and connected a network must be for computation to emerge spontaneously. A linear threshold model exhibits a percolation transition at a critical mean degree $z_{c}=1$. This means that when, on average, each node is linked to one other, a giant connected component (GCC) for nontrivial computation appears [65]. A minimal network of approximately $10^{4}$ components—if sufficiently connected and properly balanced between activation and inhibition—can begin to compute complex Boolean functions. This supports the idea that a protocell capable of decision-making may require only $10^{4}$ interacting components. If each contributes 10 bits, the total is just $10^{5}$$10^{6}$ bits—well below the earlier estimate of $10^{9}$ bits. But here, we consider only logical capacity, not structural or biochemical implementation, making this a lower bound.
Taken together, sufficiently complex chemical reaction networks can mimic the dynamics of recurrent neural networks, or make the emergence of complex logic functions inevitable. This raises the intriguing possibility that biological systems—even at the molecular scale—approach universal computation in their expressive power. Indeed, chemical kinetics, when properly structured, can implement universal computation [66]. More concretely, chemical reaction networks governed by mass-action kinetics can simulate arbitrary computations and are thus Turing complete [67].
@@ Discussion
To better understand what must be true for abiotic life to succeed, we proposed a “requirements checklist”: (i) some degree of physical or chemical bias (e.g., compartments, cycles, autocatalytic networks), (ii) sufficient persistence time if information accumulates via a biased random walk through chemical space, and (iii) protection and reuse of functional molecules. If steady progression is too slow, a phase transition may have enabled a more abrupt leap in complexity. Either way, a purely random soup is too lossy—some form of prebiotic informational structure must precede Darwinian evolution.
@@
Naturally, the framework has limitations. Our estimates—of entropy, distortion time, and protocell complexity—are crude at best, drawn from analogies rather than direct measurement. The entropy of the prebiotic chemical pool ($H_{\text{prebiotic}}$) is inferred from analogues like meteorites and Titan’s atmosphere, and the estimated information content of a minimal cell ($I_{\text{protocell}}$) is based on compressed models rather than any physical lower bound. Still, the exercise offers a quantitative scaffold for framing the problem—and perhaps guiding future investigations.
@@
Beyond parameter uncertainty, the deeper challenge is mechanistic: even if the information rate is feasible, the route remains opaque. Where did the directionality (drift $v$ or persistence $\tau$) come from? What structures or environmental constraints enabled long-term memory or error suppression without evolved proofreading? The puzzle deepens as timelines shift: from 3.465 Gy microfossils in Western Australia [15, 16, 17] to a LUCA possibly living around 4.2 Gy ago [18], close to the formation of liquid water at 4.404 Gy [13]. But LUCA is just a branching point. Substantial evolutionary development—possibly involving other, now-extinct lineages—likely occurred between the first protocell and LUCA.
@@
Which returns us, cautiously but irresistibly, to the question: Was Earth terraformed, or did order coalesce from chaos under the silent governance of physics? Today, humans seriously contemplate terraforming Mars or Venus in scientific journals [68, 69]. If advanced civilizations exist, it is not implausible they might attempt similar interventions—out of curiosity, necessity, or design. Still, Occam’s razor weighs in: abiotic evolution, however slow and strange, remains a viable (if mind-bending) explanation. Invoking terraforming adds explanatory complexity without constraint. And while we cannot prove that abiogenesis is inevitable, it remains consistent with thermodynamics.
@@
If life is the ultimate emergent phenomenon, perhaps it resists prediction not because the physics is wrong, but because the framework is incomplete. As Laughlin has argued, emergence defies reductionism not in defiance of physics, but in fulfillment of it [70, 71]. The reaction path may be immensely long, contingent, and distributed across scales—too delicate to reconstruct after four billion years. In such a landscape, quantitative feasibility may be the most we can achieve. But feasibility, even without full predictability, can still inform what is physically possible—and that, for now, is enough to keep the question alive.
@@
Looking ahead, we advocate for more empirical, exploratory work in the spirit of Miller and Urey—to better constrain $\tau$, $D$, and prebiotic entropy. We suggest focusing on mechanisms of memory and information retention in prebiotic systems. For instance, in complex and stochastic networks, do certain configurations converge to robust attractor dynamics in molecular or ecological state space? Might this make evolution more deterministic than traditionally thought, as hinted at by recent outside-of-the-box experiments [72, 73]?
@@
AI may itself play a decisive role. Beyond modeling chemical systems, it may help reverse-engineer candidate pathways, identify attractor landscapes, or uncover deep statistical regularities that escape human intuition. If life is a form of physical computation, as some have proposed [74], then AI might soon help identify how natural chemistry can become computationally powerful enough to self-organize. (Personally, I am skeptical that a generative AI can move from predicting the next word to proving the origin of life, bootstrapping us out of our existential dilemma.) Whether or not AI ultimately solves the origin-of-life problem, it may well turn out to be the next Cohen’s microscope (or telescope), just better [75]. Certainly, it has begun to rewrite the language in which biological complexity can be described.
@@
“The miracle of the appropriateness of the language of mathematics for the formulation of the laws of physics is a wonderful gift which we neither understand nor deserve.” — Eugene P. Wigner [76]
@@ Acknowledgements
We thank the Physics of Life Network of Excellence at Imperial College, for its stimulating seminars, journal club, and workshops. The latter included ”Physics of emergent behaviour III: from origin of life to multicellularity” in 2021, among others. We also acknowledge the lack of funding, which, while generally detrimental, created space for pondering the fundamental questions of this article.

### v2

@@ Abstract
An alternative possibility is that life arose via a rather sudden explosive-like transition — by some yet-unknown self-organization principle of collectiveness triumphing over reductionism, that bridged the gap between random chemistry and selective evolution, potentially circumventing the information bottleneck of early self-replication. This idea is not new. It is most well developed in the study of collectively autocatalytic sets, first introduced by Kauffman (1986) [kauffman1986autocatalytic]. In such networks, each molecule is catalyzed by another member of the set, enabling system-wide replication without requiring any single molecule to self-replicate or the presence of a membrane-enclosed protocell. This framework offers a resolution to Eigen’s paradox: long, information-rich molecules require error correction to persist, but error-correcting machinery itself demands such complexity [eigen1971self]. Autocatalytic sets enable increasing functional complexity before templated replication emerges. Mathematical models show that as chemical diversity increases, the probability of forming an autocatalytic set undergoes a sharp phase transition—a sudden onset of self-sustaining autocatalysis [hordijk2004detecting, filisetti2011stochastic]. This is analogous to a percolation transition in random graphs.
@@
Such transitions may also explain later leaps in complexity, including the emergence of the brain. Put simply, a brain functions as an internal model of the external world, trained by experience and used to predict future outcomes [hidalgo2014information] [^6] This predictive capacity has been formalized in the Bayesian brain hypothesis [Knill2004BayesianBrain], where internal priors are continuously updated in light of new data, approximating posterior beliefs that guide perception and action.. For example, one can model parameter adaptation in a toy brain using the inverse temperature $\beta$ in an Ising-like system, rather than connectivity or size, as the critical control parameter. In such models, evolutionary fitness peaks near a critical value of $\beta$ when environments are dynamic, heterogeneous, and unpredictable [hidalgo2014information]. Though phase transitions technically apply to infinite systems, the benefits of sharp transitions often hold in large finite systems as well [hidalgo2014information, depalo2017critical, mora2011criticality].
@@
In the spirit of this article, we now return to AI. Suppose that chemical reaction networks in a cell can be viewed as a kind of neural network. Then we may ask: how much information must be encoded in such a network to achieve life-like complexity? Recent work on recurrent neural chemical reaction networks (RNCRNs) has shown that artificial recurrent neural networks can be mapped onto chemical reaction networks [dack2024rncrn]. These networks are capable of approximating arbitrary dynamical systems governed by ordinary differential equations. Unlike feedforward neural networks, which approximate static input–output mappings, RNCRNs model time-dependent behaviors such as switching in multi-stable systems or oscillations frequently observed in biology.
@@
Currently, such RNCRNs are small (<10<10 species or nodes) [dack2024rncrn], and it remains unclear how large a network must be to compute the full repertoire of cellular behavior. Still, we can try to estimate the required network size. The universal approximation theorem states that for any continuous real-valued function and any target error $\varepsilon>0$, there exists a neural network that approximates $f$ within $\varepsilon$ on a compact domain. Recent results give an upper bound on the number of neurons $M$ needed to achieve this approximation:
@@
Let us now take a different route and investigate how large and connected a network must be for computation to emerge spontaneously. A linear threshold model exhibits a percolation transition at a critical mean degree $z_{c}=1$. This means that when, on average, each node is linked to one other, a giant connected component (GCC) for nontrivial computation appears [wilkerson2022spontaneous]. A minimal network of approximately $10^{4}$ components—if sufficiently connected and properly balanced between activation and inhibition—can begin to compute complex Boolean functions. This supports the idea that a protocell capable of decision-making may require only $10^{4}$ interacting components. If each contributes 10 bits, the total is just $10^{5}$$10^{6}$ bits—well below the earlier estimate of $10^{9}$ bits. But here, we consider only logical capacity, not structural or biochemical implementation, making this a lower bound.
Hence, sufficiently complex chemical reaction networks can mimic the dynamics of recurrent neural networks, or make the emergence of complex logic functions inevitable. This raises the intriguing possibility that biological systems—even at the molecular scale—approach universal computation in their expressive power. Indeed, chemical kinetics, when properly structured, can implement universal computation [magnasco1997chemical]. More concretely, chemical reaction networks governed by mass-action kinetics can simulate arbitrary computations and are thus Turing complete [soloveichik2008computation].
Taken together, these various exotic phase transitions leave the door open for the spontaneous assembly of protocells far from equilibrium. As collective emergent phenomena, such transitions are, by definition, difficult—if not impossible—to predict from mechanistic ingredients and first principles. While this may offer a route around the assembly bottleneck, it also leaves the ultimate watershed moment of life’s origin clouded in complexity-science mystery.
@@ Discussion
To better understand what must be true for abiotic life to succeed, we proposed a “requirements checklist”: (i) some degree of physical or chemical bias (e.g., compartments, cycles, autocatalytic networks), (ii) sufficient persistence time if information accumulates via a random walk through chemical space, and (iii) protection and reuse of functional molecules. If steady progression is too slow, a phase transition may have enabled a more abrupt leap in complexity. Either way, a purely random soup is too lossy—some form of prebiotic informational structure must precede Darwinian evolution.
@@
Naturally, the framework has limitations. Our estimates—of entropy, distortion time, and protocell complexity—are crude at best, drawn from analogies rather than direct measurement. The entropy of the prebiotic chemical pool ($H_{\text{prebiotic}}$) is inferred from analogues like meteorites and Titan’s atmosphere, and the estimated information content of a minimal cell ($I_{\text{protocell}}$) is based on compressed models rather than any physical lower bound. In particular, the value of factor $\eta$, likely tiny, represents a catch-all for the yet-unknown mechanism. Still, the exercise offers a quantitative scaffold for framing the problem—and perhaps guiding future investigations.
@@
Beyond parameter uncertainty, as mentioned, the deeper challenge is mechanistic: even if the information rate is feasible, the route remains opaque. Where did the directionality (drift $v$ or persistence $\tau$) come from? What structures or environmental constraints enabled long-term memory or error suppression without evolved proofreading? The puzzle deepens as timelines shift: from 3.465 Gy microfossils in Western Australia [Schopf2006, Schopf2007, Schopf2018] to a LUCA possibly living around 4.2 Gy ago [Moody2024], close to the formation of liquid water at 4.404 Gy [Wilde2001]. But LUCA is just a branching point. Substantial evolutionary development—possibly involving other, now-extinct lineages—likely occurred between the first protocell and LUCA.
@@
Which returns us, cautiously but irresistibly, to the question: Was Earth terraformed, or did order coalesce from chaos under the silent governance of physics? Today, humans seriously contemplate terraforming Mars or Venus in scientific journals [sagan1961planetary, sole2020synthetic]. If advanced civilizations exist, it is not implausible they might attempt similar interventions—out of curiosity, necessity, or design. Still, Occam’s razor weighs in: abiotic evolution, however slow and strange, remains a viable (if mind-bending) explanation. Invoking terraforming adds explanatory complexity without constraint. And while we cannot prove that abiogenesis is inevitable, it remains consistent with thermodynamics.
@@
If life is the ultimate emergent phenomenon, perhaps it resists prediction not because the physics is wrong, but because the framework is incomplete. As Laughlin has argued, emergence defies reductionism not in defiance of physics, but in fulfillment of it [Laughlin2000MiddleWay, Laughlin2005DifferentUniverse]. The reaction path may be immensely long, contingent, and distributed across scales—too delicate to reconstruct after four billion years. In such a landscape, quantitative feasibility may be the most we can achieve. But feasibility, even without full predictability, can still inform what is physically possible—and that, for now, is enough to keep the question alive.
@@
Looking ahead, we advocate for more empirical, exploratory work in the spirit of Miller and Urey—to better constrain persistence $\tau$, exploration (or drift) $D$ (or $v$), and mechanistic uncertainty $\eta$. We suggest focusing on mechanisms of memory and information retention in prebiotic systems. For instance, in complex and stochastic networks, do certain configurations converge to robust attractor dynamics in molecular or ecological state space? Might this make evolution more deterministic than traditionally thought, as hinted at by recent outside-of-the-box experiments [chuang2019homeorhesis, floroni2025membraneless]?
@@
AI may itself play a decisive role. Beyond modeling chemical systems, it may help reverse-engineer candidate pathways, identify attractor landscapes, or uncover deep statistical regularities that escape human intuition. If life is a form of physical computation, as some have proposed [Lloyd2000], then AI might soon help identify how natural chemistry can become computationally powerful enough to self-organize.
Although current AI systems excel at synthesizing knowledge and exploring enormous hypothesis spaces, their creativity remains largely one of rearrangement rather than genuine conceptual invention. I am personally sceptical that deep scientific insight can be achieved without the qualities that shape human creativity — purpose, passion, embodiment, and an awareness of our own finiteness. These give humans a sense of what truly matters and drive the irrational persistence behind transformative ideas. For instance, a lone genius might have a breakthrough driven by deeply personal, often irrational motivations  [^7] History offers countless examples in which creative persistence was fueled by romantic, social, or symbolic incentives rather than rational optimization. That kind of urgency or drive is hard to explain to an AI.. Still, future AI may either approximate such motivational structures or overcome their absence through overwhelming exploratory capacity. For now, however, the kind of abductive leaps required to explain life’s emergence remain uniquely human. Whether or not AI ultimately solves the origin-of-life problem, it may nevertheless turn out to be the next Cohen’s microscope (or telescope), just better [cohen2004mathematics]. Certainly, it has begun to rewrite the language in which biological complexity can be described.
As a thought experiment, one can ask what a contemporary large language model might propose if tasked with addressing the origin-of-life problem under effectively unlimited resources over a decade. The model proposes a surprisingly coherent strategy. First, build a digital twin of early Earth—and of selected exoplanets—to assess whether our planet’s conditions were in any sense special. Second, create a vast “chemistry collider” capable of exploring prebiotic reactions at scale, in order to map the chemical landscape and understand the striking absence of mid-complexity abiotic molecules today. Third, systematically construct minimal forms of artificial life to determine what is truly required for sustained heredity, metabolism, and evolvability. Finally, establish a serious theory institute: a place where a few deeply creative individuals (“lone geniuses”), largely unconstrained by the burdens of everyday academic life, can think deeply about the problem, guided by the data above, and develop theoretical frameworks that go beyond purely brute-force approaches.
The striking point is that neither brute-force exploration nor isolated genius is likely to solve the problem on its own. Progress will almost certainly emerge from the interplay—perhaps the tension (likely frustration) — between large-scale big data with even bigger computers and small-scale conceptual innovation. Even so, we may never recover the precise historical sequence by which life first emerged. What we can hope for, however, is a set of physically and chemically plausible routes: an understanding of how matter of this kind could have become alive, even if we never know exactly how it did.
@@
“The miracle of the appropriateness of the language of mathematics for the formulation of the laws of physics is a wonderful gift which we neither understand nor deserve.” — Eugene P. Wigner [Wigner1960]
Data and Code Availability
No experimental data was produced in this work. All original code can be found in the GitHub repository https://github.com/Endres-group/origin-of-life-paper and is publicly available.
@@ Acknowledgements
We thank the Physics of Life Network of Excellence at Imperial College, for its stimulating seminars, journal club, and workshops. The latter included ”Physics of emergent behaviour III: from origin of life to multicellularity” in 2021, among others. We also appreciate the hospitality at the Nature conference ”AI augmented Biology” in Nanjing, China, in October 2025, which led to stimulating discussions. We also acknowledge the lack of funding, which, while generally detrimental, created space for pondering the fundamental questions of this article.
Summary of changes in Version 2
Version 2 expands the conceptual scope of the paper by discussing the role of the global magnetic field, clarifying different photosynthetic pathways and their implications for the lifetime of the biosphere, and using the homochirality problem as an illustrative contribution to the efficiency factor $\eta$. It extends the protocell information estimate to include assembly-path complexity, refines the analysis of persistence times in random-walk-based assembly (with new Fig. 2B), introduces an unbalanced optimal transport bound on the minimal energetic cost of protocell formation, and adds a new section on assembly theory supported by a toy model (Fig. 3). The Discussion has been broadened to include a more critical perspective on the potential and limitations of artificial intelligence in origin-of-life research. Finally, a GitHub repository has been added and several references have been corrected.